110 likes | 200 Views
Unstable vref for 1 sift on board 667 (1B7). Board was removed and inspected by M. Matulik, nothing wrong was found on the board, board was re-inserted. Vref change: last_calibration-previous_calibration. Unstable vref for 1 sift on board 667 (1B7).
E N D
Unstable vref for 1 sift on board 667 (1B7) • Board was removed and inspected by M. Matulik, nothing wrong was found on the board, board was re-inserted. • Vref change: last_calibration-previous_calibration
Unstable vref for 1 sift on board 667 (1B7) • Unstable vref causes spike on hot map for fps
Unstable vthres for board 557 (6A5) • Hot region developed on 3/11:
Unstable vthres for board 557 (6A5) • Board was re-calibrated on 3/11, change of vthres with repect to the previous calibration: • Board stable since 3/11
New "latched" VLPC, 3B1 • Now 7 "latched" VLPCs in the system.
Non working channels, looking for groups to identify bad VLPCs, sifts
Non working channels in CFTX3 • Fibers 15-18 and 47-50 are on 6A0, SVX 7 all readout by VLPC 8; board was already swapped, so the problem is inside cassette.
Non working channels in cft according e-log on 3/4/2003 and 3/13/2003 • CFTX2: 9 fibers between 1347 and 1406, 8 belong to VLPC7 in 9A5 svx0 • CFTX3: 11 fibers between 15 and 50, 8 belong to VLPC8 in 6A0 svx7 • This board was swapped and the problem remains-cassette problem • CFTX5: 11 fibers between 2183 and 2233, 8B6 MCM7 should be swapped • CFTX6: 9 fibers between 1167 and 1202, 8 belong to VLPC8 in 5A4 svx7 • CFTX6: 9 fibers between 4611 and 4670, 8 belong to VLPC2 in 7A2 svx7 • Not on list of unbiased VLPCs • CFTX7: 9 fibers beween 909 and 948, 8 belong to VLPC7 in 6A2 svx7 • CFTX8: 10 fibers between 4237 and 4276, 8 belong to VLPC 7 in 8A6 svx0, NEWONLY ON 3/13/2003 ! Not on list of unbiased VLPCs • CFTU3: 8 fibers between 4301 and 4340, all belong to VLPC7 in AA6 svx0 • Not on list of unbiased VLPCs • CFTU4: 8 fibers between 5412 and 5489, 8B2 and 8B3
Problems with board 637 • Board 637 in slot 3A7 • Looses bias read-back, needs to be power-cycled to recover bias read-back3A7 board sometimes stops working. We had recently three instances: 2/19/2993 eLog 177223 2/26/2003 eLog 180308 3/13/2003 eLog 184123
Problems with power suppliesMajor alarm form CFT_LVAFE_PS2-B/+55
Problems with power supplies • FPD power supply died over the weekend • Disabled from alarm system: • CFT_LVAFE_PSC-A/+12V • eLog 2/13/2003: readback error • CFT_LVAFE_PS2-B/+33V